OpenStudy (anonymous):

3y^2+5Y+2 does anybody know how to factor this one?>

OpenStudy (unklerhaukus):

\[ay^2+by+c=0\] use the quadratic formula \[y_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[(y-y_1)(y-y_2)=0\]

OpenStudy (phoenixfire):

(y+1)(y+2/3) It's the same as your last question. Use the quadratic formula to get your y values.